Terrell is thinking of 2 numbers. The sum of the numbers is 9. One of the numbers is 3 less than twice the other number. Which of the numbers below is the greater number?
The choices are: 6, 5, 4, 3 or 2

Answer by ad_alta(240)

Let 'n' be the greater number, and 'm the smaller. Then n+m=9 and 2m=n+3. Substituting the first equation into the second we get 2m=(9-m)+3, so m=4 and n=5. Thus the greater number is 5.